# Optimally Deep Networks - Adapting Model Depth to Datasets for Superior Efficiency

### v1

@@ III-D Fine-tuning the Optimally Deep Network
Once the Optimally Deep Network is obtained, it is fine-tuned till full convergence to maximize its performance as shown in Figure 2(c). The fine-tuning stage is the final step in training Optimally Deep Networks (ODNs) to ensure the selected optimal depth achieves its best possible performance. Fine-tuning allows the model parameters with optimal depth to adjust more precisely to the target dataset, further improving accuracy (without altering the depth). Afterwards, the ODN can be extracted (by copying weights) for deployment.
@@ IV Experiments & Results
We evaluate Optimally Deep Networks (ODNs) across five benchmark datasets of varying complexity: MNIST, EMNIST, Fashion-MNIST, SVHN, and CIFAR-10. All the datasets have 10 classes except EMNIST, which has 26 classes (for “letters”). To demonstrate the simplicity and applicability of our approach, we use three widely adopted ResNet architectures: ResNet-18, ResNet-34, and ResNet-50. We train the models using SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 128, and initial learning rate of 0.1. The code of Optimally Deep Networks (ODNs) is available at: https://github.com/saktx/Optimally_Deep_Networks.
@@
Table I presents the performance of ODNs as compared to networks with full depths. It can be observed that the optimally deep ResNet-18 trained on MNIST and EMNIST achieves a comparable accuracy of 99.31 % and 95.04 % while reducing the model size (memory footprint) by 98.64 % and 96.54 %, respectively. Only 2 out of the 8 blocks of ResNet-18 are sufficient to match the complexity of the MNIST dataset whereas only 3 blocks are sufficient for EMNIST. Similarly, the optimally deep ResNet-34 achieves accuracies of 93.54 % and 96.08 %, with model size reductions of 95.04 % and 96.44 % on Fashion-MNIST and SVHN, respectively. For the CIFAR-10 dataset, the optimally deep ResNet-50 with a depth of 11 blocks instead of 16, provides the accuracy of 93.35 % with a reduction of 73.06 % in model size. Compared to full-depth models, the accuracies of ODNs remain within a 1.75 % tolerance, while the memory footprint reductions are substantial, exceeding 95 % for low-complexity datasets: MNIST, EMNIST, Fashion-MNIST, and SVHN.
Table I also shows that the substantial reduction in parameter count and model size of ODNs leads to a decrease in FLOPs. Lower FLOPs result in faster inference times and reduced latency in edge devices. For MNIST and EMNIST, the memory footprint is reduced from 44.78 MB to 0.61 MB and 1.55 MB, respectively. Similarly, for Fashion-MNIST and SVHN, the memory footprint decreases from 85.29 MB to 4.23 MB and 3.04 MB, respectively. In the case of CIFAR-10, the memory footprint is reduced by 73.06 %, from 94.43 MB to 25.44 MB, reflecting the higher complexity of this dataset.
ODNs eliminate the need for an expensive and complex search space as compared to complex NAS methods. Unlike sparsification and pruning, ODNs achieve a substantial reduction in the memory footprints of models. In contrast to Once-for-All training, ODNs are simpler, requiring only a single progressively expanding model without the complexity of maintaining a super-net [22]. Another advantage of ODNs is that in order to find the optimal depth, the model is trained till convergence using shallower depths during progressive depth expansion stage. These fully converged intermediate models can be saved as by-products at each depth and later used for deployment on devices when stricter accuracy–efficiency trade-offs are required.
@@ Conclusion
We propose Optimally Deep Networks (ODNs), models which adapt their depth according to the complexity of the task at hand. Typically, powerful architectures are trained at full depths but not all datasets or tasks require such high model capacity. ODNs are trained using progressive depth expansion: training begins with shallow depths and the network is incrementally deepened only as needed, until the desired performance is achieved. This approach enables the extraction of optimally deep networks that significantly reduce memory footprint, computational overhead, and inference cost, while maintaining competitive accuracy. Empirical evaluations on MNIST, EMNIST, Fashion-MNIST, SVHN, and CIFAR-10 using ResNet-18, ResNet-34, and ResNet-50 architectures demonstrate that ODNs can achieve up to 98.64 % reduction in model size for low-complexity datasets while maintaining good performance. ODNs present a practical framework for searching optimally deep models that balance resource efficiency with accuracy, paving the way for scalable and deployable neural networks in real-world applications.

### v4

@@ III-D Fine-tuning the Optimally Deep Network
Once the Optimally Deep Network is obtained, it is fine-tuned till convergence to maximize its performance as shown in Figure 3(c). The fine-tuning stage is the final step in training Optimally Deep Networks (ODNs) to ensure the selected optimal depth achieves its best possible performance. Fine-tuning allows the model parameters with optimal depth to adjust more precisely to the target dataset, further improving accuracy (without altering the depth). The ODN can be extracted from the backbone at the end (by copying weights) for deployment.
@@ IV Experiments & Results
We evaluate Optimally Deep Networks (ODNs) across five benchmark datasets of varying complexity: MNIST, EMNIST, SVHN, Fashion-MNIST, and CIFAR-10. All the datasets have 10 classes except EMNIST, which has 26 classes (for “letters”). To demonstrate the simplicity and applicability of our approach, we use three widely used architectures: ResNet-18, ResNet-34, and ResNet-50. We train the models using SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 128, and initial learning rate of 0.1.
@@
Table I presents the performance of ODNs as compared to networks with full depths. It can be observed that the optimally deep ResNet-18 trained on MNIST and EMNIST achieves a comparable accuracy of 99.31 % and 95.04 % while reducing the model size (memory footprint) by 98.64 % and 96.54 %, respectively. Only 2 out of the 8 blocks of ResNet-18 are sufficient to match the complexity of the MNIST dataset whereas only 3 blocks are sufficient for EMNIST. Similarly, the optimally deep ResNet-34 achieves accuracies of 96.08 % and 93.54 %, with model size reductions of 96.44 % and 95.04 % on SVHN and Fashion-MNIST, respectively. For the CIFAR-10 dataset, the optimally deep ResNet-50 with a depth of 11 blocks instead of 16, provides the accuracy of 94.12 % with a reduction of 73.06 % in model size. Compared to full-depth models, the accuracies of ODNs can be achieved within 1% tolerance, while the memory footprint reductions are substantial, exceeding 95 % for low-complexity datasets: MNIST, EMNIST, SVHN, and Fashion-MNIST.
Table I also shows that the substantial reduction in parameter count and model size of ODNs leads to a decrease in FLOPs. Lower FLOPs result in faster inference times and reduced latency in edge devices. For MNIST and EMNIST, the memory footprint is reduced from 44.78 MB to 0.61 MB and 1.55 MB, respectively. Similarly, for SVHN and Fashion-MNIST, the memory footprint decreases from 85.29 MB to 3.04 MB and 4.23 MB, respectively. In the case of CIFAR-10, the memory footprint is reduced by 73.06 %, from 94.43 MB to 25.44 MB, reflecting the higher complexity of this dataset.
ODNs eliminate the need for an expensive and complex search space as compared to complex NAS methods. Unlike sparsification and pruning, ODNs achieve a substantial reduction in the memory footprints of the models. In contrast to Once-for-All (OFA) training, ODNs are simpler, requiring only a single progressively expanding model without the complexity of maintaining a super-net [OFA_1]. Another advantage of ODNs is that in order to find the optimal depth, the model is trained till convergence using shallower depths during progressive depth expansion stage. These fully converged intermediate models can be saved as by-products at each depth and later used for deployment on devices (like OFA) when stricter accuracy-efficiency trade-offs are required.
@@ Conclusion
We propose Optimally Deep Networks (ODNs), models which adapt their depth according to the complexity of the task at hand. Typically, powerful architectures are trained at full depths but we show that not all datasets require high model capacity. ODNs are trained using progressive depth expansion: training begins with shallow depths and the network is incrementally deepened only as needed, until the desired performance is achieved. This enables the extraction of Optimally Deep Networks that significantly reduce memory footprint, computational overhead, and inference cost, while maintaining competitive accuracy. Empirical evaluations on MNIST, EMNIST, SVHN, Fashion-MNIST, and CIFAR-10 using ResNet-18, ResNet-34, and ResNet-50 show that ODNs can achieve up to 98.64 % reduction in model size while maintaining good performance. ODNs present a practical framework for searching optimally deep networks that balance resource efficiency with accuracy, paving the way for scalable and deployable neural networks in real-world applications.